f(y)=sinx+cscx²y+2xy
1
Expert's answer
2020-02-04T10:53:55-0500

"f(x,y)=sin(x)+csc(x^2y)+2xy"

"\frac{\partial f}{\partial x}=cos(x)-2xy\frac{cos(x^2y)}{sin^2(x^2y)}+2y"

"\frac{\partial f}{\partial y}=-x^2\frac{cos(x^2y)}{sin^2(x^2y)}+2x"